1.  If 15.0 L Neon at 60 0C is heated at constant pressure, what temperature will it have if it occupies a volume of 4.50 L?


1
Expert's answer
2022-05-03T09:03:05-0400

Given:

V1 = 15.0 L

T1 = 60°C + 273.15 = 333.15 K

P1 = P2 = constant

V2 = 4.50 L

T2 = ???


Solution:

Since the pressure and amount of gas remain unchanged, Charles's law can be used.

Charles's law can be expressed as: V1 / T1 = V2 / T2

V1T2 = V2T1

To find the temperture, solve the equation for T2:

T2 = V2T1 / V1

T2 = (4.50 L × 333.15 K) / (15.0 L) = 99.945 K = 100 K

T2 = 100 K or −173°C


Answer: at 100 K